Please help me solve this problem.
If the perimeter of a rectangular closet is 40 feet and the width is one foot less than half the length, what are the dimensions of the length and width?

Answer by thejackal(72)

If P = 2L + 2W where L and W are Length and Width Respectively
then
40 = 2L + 2W;
since W = (L/2)-1
then 40 = 2L + 2((L/2)-1)
simplified to 40 = 3L - 2
42/3 = L
solve for L you will find that L = 14
plug that into the equation for W such
that
W = (14/2)-1 = 6